(1 pt)

Solve the given system, using the substitution method.

y = 3x – 4
9x – 3y = 14




A.

There is no solution.


B.

(12, 14)


C.

(14, 12)


D.

There are an infinite number of solutions.

Answer by stanbon(75887)

Solve the given system, using the substitution method.
y = 3x – 4
9x – 3y = 14
----
9x-3(3x-4) = 14
-----
9x - 9x + 12 = 14
---
12 = 14
-------
That is a contradiction.
Ans: No solution
-------------------
Cheers,
Stan H.